# The Relative Entropy of Expectation and Price

### v4

@@ Abstract
for the price $p$ and payoff $P$ of a funded and hedged derivative security, where the price measure $\mathbb{E}$ has minimum entropy relative to economic expectations, and the parameter $\alpha$ matches the risk aversion of the investor. This price principle is easily applied to standard models for market evolution, with applications considered here including model risk analysis, deep hedging and decentralised finance.
@@ I Introduction
A derivative transaction exchanges the initial price $p$ for a commitment to return the contractual final price $p+dp$ at a later time. In determining its economic viability, the counterparties assess whether the trade enhances or degrades the overall performance of their investment portfolios, balancing expected return against net risk. Asymmetry between the expectations and investment objectives of participants then drives market evolution, following the dynamic that trades execute when parties agree on price but disagree on value.
@@
Incentivising margin
In this approach, the investment strategy accounts for both the funding costs and risk management benefits of margin, and quantifies this in the derivative valuation. Margin thus becomes a key consideration for pricing.
@@
Uniquely identifying price from expectation, market and model risks are distinguished respectively as sensitivities to observed prices and subjective economic assumptions. This enables a hierarchical approach that recognises market tiering.
@@
Advocates of deep hedging replace risk neutralisation with risk optimisation as the guiding principle for pricing, as data-driven models see the complex correlations among economic variables as opportunity rather than hindrance. Convex risk optimisation is similarly used here to derive the price measure from the expectation measure, though the objectives differ. Model risk is quantified as the residual risk after funding and hedging, which requires for its definition a precise understanding of the relationship between expectation and price. Taking this idea to its logical conclusion, the convex risk metric is implemented as margin held against model risk. Optimisation of the strategic return net of all settlements then naturally embeds risk optimisation.
@@
Suppose that two counterparties deploy initial capital $c_{1}>0$ and $c_{2}>0$ to the transaction and deposit the proceeds in a liquid margin account with positive unit price $b>0$. The first counterparty buys the derivative from the second counterparty, and the capital deployed must be sufficient to enter the transaction:
@@
With the proceeds and unused capital deposited in the margin account, the final margin available to settle the derivative is $(c_{1}-p)(1+db/b)$ for the first counterparty and $(c_{2}+p)(1+db/b)$ for the second counterparty. This is sufficient to settle the contractual commitments of the derivative when:
$-(c_{1}-p)(1+db/b)\leq p+dp\leq(c_{2}+p)(1+db/b)$ (3)
@@
This simple framework for bilateral margin is used in the following to derive equilibrium pricing expressions that account for funding and default. It is an idealisation of real trading activity, and in particular does not account for the default protections provided by additional capital covering the net risks of the institution. Extensions to multilateral margin, which can be incorporated into the framework, are not considered here.
In this essay, foundations for pricing are considered that do not rely on unrealistic assumptions on market completeness or the effectiveness of the trading strategy, properly account for residual model risk, and establish consistency across diverse applications including algorithmic trading, derivative pricing and hedging, and the modelling of margin and capital. In developing these foundations, the entropic risk metric has dual application: as the a priori measure of performance used to optimise the investment strategy; and as the determinant of margin used to protect against default.
@@ II Entropic investment strategies
The objective of the investor is to optimise the return $dc$ on the investment strategy, a task that depends on an initial assessment of the final economy that is both subjective and indeterminate. This assessment is encapsulated in the expectation measure $\mathbb{E}$ of the investor. Define the corresponding probability, variance and covariance operators:
$\mathbb{P}[\mathcal{A}]$ $:=\mathbb{E}[1_{\mathcal{A}}]$ (8)
$\mathbb{V}[X]$ $:=\mathbb{E}[XX^{t}]-\mathbb{E}[X]\mathbb{E}[X]^{t}$
@@
for the event $\mathcal{A}$ and the random vectors $X$ and $Y$. Investment performance is estimated using these operators, and the strategy is optimised accordingly.
@@ II.1 Entropic risk metric
where $\mathbb{M}_{n}$ is the $n$th cumulant operator of the investor, so that $\mathbb{M}_{0}=0$, $\mathbb{M}_{1}=\mathbb{E}$ and $\mathbb{M}_{2}=\mathbb{V}$. The change of measure implicit in this definition reweights according to the sign and magnitude of the return, with properties that make it a useful metric of value-at-risk.
@@
is monotonic decreasing as a function of the parameter $\alpha>0$ and has limits:
$\mathbb{E}[0][dc]$ $=\mathbb{E}[dc]$ (12)
@@
Instances of this equality for $\bar{\mathbb{E}}=\mathbb{E}$ and $\bar{\mathbb{E}}=\mathbb{E}^{\alpha}$ generate bounds for the entropy-adjusted mean:
[…]
@@
These are standard results from information theory, and only the key elements of their proofs are presented here. Monotonicity is a consequence of Hölder’s inequality:
@@
for $0<\alpha_{1}<\alpha_{2}$, so that $\mathbb{E}[\alpha_{1}][dc]\geq\mathbb{E}[\alpha_{2}][dc]$. The first limit follows from the expansion of the cumulant generating function. For the second limit, Markov’s inequality:
@@
The lever $\alpha>0$ thus dials the entropy-adjusted mean $\mathbb{E}[\alpha][dc]$ across the span of possible values for $dc$ between $\operatorname*{ess\,inf}[dc]$ and $\mathbb{E}[dc]$ and no further. Entropic risk is, in this sense, more targeted than variance adjustment as a moments-based metric of value-at-risk. The equivalent measure $\mathbb{E}^{\alpha}$ plays a central role in this result, and the gap between the mean $\mathbb{E}[dc]$ and the entropy-adjusted mean $\mathbb{E}[\alpha][dc]$ is proportional to the distance between the measure $\mathbb{E}$ and the measure $\mathbb{E}^{\alpha}$, as defined by their relative entropy.
@@ II.2 Optimising entropic risk
The investor has a choice of market strategies whose returns over the investment period are the components of the return vector $dq$. The return on the investment portfolio whose weights are the components of the weight vector $\omega$ is then $dc:=\omega\cdot dq$, the weighted sum of the returns on the market strategies.
@@
Risk aversion is controlled by the parameter $\alpha$ in the range $0\leq\alpha\leq\infty$, and entropic risk is monotonic in this range:
@@
$\phi^{(0)}$ $=0$ (25)
$\phi^{(n+1)}$ $=\phi^{(n)}+\mathbb{V}^{\phi^{(n)}}[dq]^{-1}\mathbb{E}^{\phi^{(n)}}[dq]$
calibrates the unit optimal portfolio to the calibration condition when the variance is invertible. This portfolio is a property of the market return, generating the price measure $\mathbb{E}^{\phi}$ equivalent to the expectation measure $\mathbb{E}$.
@@
where the calibration conditions in the $i$th row maximise the entropy-adjusted mean of the $i$th optimal portfolio $\phi^{i}=\phi_{1}^{i}\oplus\cdots\oplus\phi_{i}^{i}$ constituted in the lowest $i$ market tiers only. The $i$th investor then takes responsibility for determining the primary component $\psi_{i}$ of the $i$th optimal portfolio and its delta $\delta^{i}=\delta_{1}^{i}\oplus\cdots\oplus\delta_{i-1}^{i}$ to the $(i-1)$th optimal portfolio in the underlying sub-market. Stacking these deltas for $i=1,\ldots,n$ generates the optimal portfolio of the full market.
Theorem II.2 (Optimising entropic risk).
Consider a market of investment strategies whose return vector decomposes as a direct sum $dq=dq_{1}\oplus\cdots\oplus dq_{n}$ where $dq_{i}$ is the return vector for the $i$th tier of the market. For each $i=1,\ldots,n$, define the $i$th sub-market with return vector $dq^{i}=dq_{1}\oplus\cdots\oplus dq_{i}$ and let $\phi^{i}=\phi_{1}^{i}\oplus\cdots\oplus\phi_{i}^{i}$ be a portfolio in this sub-market. Use the invertible linear change of variables:
@@
to represent the delta from the portfolio $\phi^{i-1}$ to the portfolio $\phi^{i}$ as the primary portfolio $\psi_{i}$ in the leading market tier minus the hedge portfolio $\delta_{1}^{i}\oplus\cdots\oplus\delta_{i-1}^{i}$ in the underlying market tiers.
Let $\mathbb{E}$ be a probability measure. For each $i=0,\ldots,n$, define the equivalent probability measure $\mathbb{E}^{i}$ with Radon-Nikodym weight:
@@
for $j=1,\ldots,i$. Starting with the probability measure $\mathbb{E}^{0}=\mathbb{E}$, these calibrations are iterated to obtain the unit optimal portfolio $\phi^{n}$ in the full market, with its associated equivalent probability measure $\mathbb{E}^{n}$ calibrated to the calibration conditions for the market strategies.
Proof.
This result follows from the measure-change relationship:
$\log{}$ $\mathbb{E}^{i-1}\exp[-(\psi_{i}\cdot dq_{i}-(\delta_{1}^{i}\cdot dq_{1}+\cdots+\delta_{i-1}^{i}\cdot dq_{i-1}))]$
[…]
$-\log\mathbb{E}\exp[-(\phi_{1}^{i-1}\cdot dq_{1}+\cdots+\phi_{i-1}^{i-1}\cdot dq_{i-1})]$
between the entropy-adjusted means. ∎
@@ III Entropic pricing and hedging
The market of underlying securities accessible to the investor performs the two key functions of funding and hedging for the derivative security. Fix a finite investment horizon, and assume the market is tiered into a set of underlying securities with initial price vector $q$ and final price vector $q+dq$ and a derivative security with initial price $p$ and final price $p+dp$. The investor quantifies a priori the performance of the investment strategy using an initial assessment of the final economic state, encapsulated in the expectation measure $\mathbb{E}$ that maps variables at final time $t+dt$ to their expectations at initial time $t$. Following the strategy developed in the previous section, the optimal portfolio then maximises the entropy-adjusted mean return at a specified risk aversion $\alpha$ for the investor.
@@
where $(r+\alpha s)\,dt$ is the Lagrange multiplier used to implement the self-funding condition $\delta\cdot q=p$.
@@
Entropic Pricing and Hedging
Fix a finite investment horizon and let $\mathbb{E}$ be the expectation measure mapping variables at the final time $t+dt$ to their expectations at the initial time $t$. Assume the market is tiered into a set of underlying securities with initial price vector $q$ and final price vector $q+dq$ and a derivative security with initial price $p$ and final price $p+dp$. Consider two self-funded strategies: one comprising underlying securities only; and one that also includes the derivative security.
These two strategies offer different entropy-adjusted mean returns to the investor. When the derivative security has no intrinsic value beyond its contractual link with the underlying securities, the efficient market equilibrates to the indifferent state wherein these entropy-adjusted means align, leading to the price condition: $\mathbb{E}[\alpha][\omega\cdot dq]=\mathbb{E}[\alpha][dp+(\omega-\delta)\cdot dq]$ (42) The calibration condition is solved to derive the optimal portfolio $\omega$ of underlying securities. The hedge and price conditions are then solved simultaneously to derive the initial price $p$ from the final price $p+dp$ of the derivative security and the hedge portfolio $\delta$ of underlying securities that re-establishes optimality for the combined portfolio.
Refer to caption Refer to caption
$\rho=-0.5$ $\rho=-0.5$
Refer to caption Refer to caption
$\rho=0$ $\rho=0$
Refer to caption Refer to caption
$\rho=0.5$ $\rho=0.5$
@@
Figure 6: Discounting is an emergent feature in entropic pricing, where the funding rate is implied as the Lagrange multiplier that implements the self-funding condition. In this example, two underlying securities are modelled with lognormal final prices $q+dq=q(1+\mu\,dt)\exp[\sigma\sqrt{dt}\,z-\sigma^{2}dt/2]$ in the expectation measure, using standard normal variables $z$ that have correlation $\rho$. The model for the second underlying security is fixed with $q_{2}=1$, $\mu_{2}=5\%$ and $\sigma_{2}=20\%$. The model for the first underlying security uses $q_{1}=1$, and the graphs in each row plot the implied funding rate $r$ and unit optimal portfolio $\phi_{1}$ as $\mu_{1}$ and $\sigma_{1}$ are varied. The correlation is $\rho=-0.5$ in the first row, $\rho=0$ in the second row, and $\rho=0.5$ in the third row, and the investment horizon is set to $dt=1$.
      When the volatility $\sigma_{1}$ of the first underlying security is small, the implied funding rate $r$ aligns with the mean rate $\mu_{1}$, as in this case the underlying security approximates a funding security with predictable return. When the volatilities $\sigma_{1}$ and $\sigma_{2}$ of the underlying securities are close to each other, the implied funding rate $r$ is approximately the average of the two mean rates $\mu_{1}$ and $\mu_{2}$. The mean rates of the underlying securities thus contribute to the implied funding rate, which is essentially the predictability-weighted average of the mean rates.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 7: Three examples of entropic hedging and pricing for a call option on an underlying security with return $dq$ modelled as a normal variable, a Poisson variable with ten expected jumps, and a uniform variable in the price measure. In each case, the return is standardised by linear transformation to have zero mean and unit standard deviation, and the derivative payoff for strike $k$ is $P:=(dq-k)^{+}$. Risk aversion is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$. The graphs plot the delta and price per unit notional as a function of the strike for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Residual risk after hedging creates a bid-offer spread around the mid price and skews the optimal hedge ratio for positions with large notional.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 8: Three examples of entropic hedging and pricing for a digital option on an underlying security with return $dq$ modelled as a normal variable, a Poisson variable with ten expected jumps, and a uniform variable in the price measure. In each case, the return is standardised by linear transformation to have zero mean and unit standard deviation, and the derivative payoff for strike $k$ is $P:=1_{dq>k}$. Risk aversion is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$. The graphs plot the delta and price per unit notional as a function of the strike for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Residual risk after hedging creates a bid-offer spread around the mid price and skews the optimal hedge ratio for positions with large notional.

### v5

@@ Abstract
for the price $p$ and payoff $P$ of a funded and hedged derivative security, where the price measure $\mathbb{E}$ has minimum entropy relative to economic expectations, and the parameter $\alpha$ matches the risk aversion of the investor. This price principle is easily applied to standard models for market evolution, with applications considered here in model risk analysis, deep hedging and quantum information.
@@ I Introduction
A derivative transaction exchanges the execution price $p$ at initial time $t$ for a commitment to return the contractual price $p+dp$ at final time $t+dt$. In determining its economic viability, the counterparties assess whether the trade enhances or degrades the overall performance of their investment portfolios, balancing expected return against net risk. Asymmetry between the expectations and investment objectives of participants then drives market evolution, following the dynamic that trades execute when parties agree on price but disagree on value.
@@
Information model
Entropic risk optimisation is applicable wherever the concept of entropy is defined. This includes theoretical or data-driven models with classical or quantum information. Entropic pricing unlocks the novel algorithmic potential of these frameworks.
@@
Uniquely determining price from expectation, market and model risks are distinguished respectively as sensitivities to observed prices and subjective economic assumptions. This enables a hierarchical approach that recognises market tiering.
@@
Advocates of deep hedging replace risk neutralisation with risk optimisation as the guiding principle for pricing, as data-driven models see the complex correlations among economic variables as opportunity rather than hindrance. Risk optimisation is similarly used here to derive the price measure from the expectation measure, though the objectives differ. Model risk is quantified as the residual risk after funding and hedging, which requires for its definition a precise understanding of the relationship between expectation and price. Taking this idea to its logical conclusion, the convex risk metric is implemented as margin held against model risk. Optimisation of the strategic return net of all settlements then naturally embeds risk optimisation.
@@
Suppose that two counterparties deploy initial capital $c_{1}>0$ and $c_{2}>0$ to the transaction and deposit the proceeds in a liquid margin account with positive unit price $b>0$. The first counterparty buys the derivative from the second counterparty, and the capital must be sufficient to enter the transaction:
@@
The capital $c_{1}$ deployed by the first counterparty is partitioned into the variation margin $p$ and the initial margin $m_{1}:=c_{1}-p$, and the capital $c_{2}$ deployed by the second counterparty is partitioned into the variation margin $-p$ and the initial margin $m_{2}:=c_{2}+p$. With the initial margin deposited in the margin account, the final margin available to settle the derivative is $m_{1}(1+db/b)$ for the first counterparty and $m_{2}(1+db/b)$ for the second counterparty. This is sufficient to settle the contractual commitments of the derivative when:
$-m_{1}(1+db/b)\leq p+dp\leq m_{2}(1+db/b)$ (3)
@@
Fixing a target for the investment strategy, the expected performance is quantified by the entropy-adjusted mean return net of funding and default:
where the measure $\mathbb{E}$ encapsulates the expectations of the investor and the parameter $\alpha>0$ controls their risk aversion. In this strategy, the trade is considered to be viable when the entropy-adjusted mean is positive. Accounting for the expectations and risk appetites of both counterparties, the trade is viable when:
$\mathbb{E}_{2}[-\alpha_{2}][-dc_{2}]\leq 0\leq\mathbb{E}_{1}[\alpha_{1}][dc_{1}]$ (9)
Setting both sides of this window of viability to zero imposes two fair pricing conditions on the initial and variation margins. The remaining degree of freedom controls the level of default protection embedded in the trade.
This simple framework for bilateral margin is used in the following to derive equilibrium pricing expressions that account for funding and default. It is an idealisation of real trading activity, and does not account for the default protections provided by additional capital covering the net risks of the institution. Extensions to multilateral margin, which can be incorporated into the framework, are not considered here.
In this essay, entropic foundations for pricing are investigated that do not rely on unrealistic assumptions regarding market completeness or the effectiveness of the trading strategy, properly account for residual model risk, and establish consistency across diverse applications including algorithmic trading, derivative pricing and hedging, and the modelling of margin and capital. In developing these foundations, the entropic risk metric has dual application: as the a priori measure of performance used to optimise the investment strategy; and as the determinant of margin used to protect against default. These results leverage the remarkable versatility and tractability of relative entropy as a universal metric of uncertainty, applicable across a wide range of models and algorithms.
@@ II Entropic investment strategies
The objective of the investor is to optimise the return $dc$ on the investment strategy, a task that depends on an initial assessment of the final economy that is both subjective and indeterminate. This assessment is encapsulated in the expectation measure $\mathbb{E}$ of the investor, which maps the final variable $X$ to its initial expectation $\mathbb{E}[X]$. Use the following notation for the associated operators of variance, covariance and probability:
$\mathbb{V}[X]$ $:=\mathbb{E}[XX^{t}]-\mathbb{E}[X]\mathbb{E}[X]^{t}$ (10)
@@
$\mathbb{P}[\mathcal{A}]$ $:=\mathbb{E}[1_{\mathcal{A}}]$
for the random vectors $X$ and $Y$ and the event $\mathcal{A}$, where $1_{\mathcal{A}}$ is the indicator of the event. Investment performance is estimated using these operators, and the strategy is optimised accordingly.
@@ II.1 Entropic risk metric
where $\mathbb{M}_{n}$ is the $n$th cumulant operator of the investor, with $\mathbb{M}_{0}=0$, $\mathbb{M}_{1}=\mathbb{E}$ and $\mathbb{M}_{2}=\mathbb{V}$. The change of measure implicit in this definition reweights according to the sign and magnitude of the return, with properties that make it a useful metric of value-at-risk.
@@
satisfies the linearity property:
$\mathbb{E}[\alpha][\lambda\,dc+\mu]=\lambda\mathbb{E}[\lambda\alpha][dc]+\mu$ (14)
for the scalars $\lambda$ and $\mu$ and is monotonic decreasing as a function of the adjustment parameter $\alpha$ with limits:
$\mathbb{E}[-\infty][dc]$ $=\operatorname*{ess\,sup}[dc]$ (15)
$\mathbb{E}[0][dc]$ $=\mathbb{E}[dc]$
@@
Linearity follows directly from the definition of the entropy-adjusted mean. Without loss of generality, assume the adjustment parameter is positive and use linearity to derive the corresponding results for the case of negative adjustment parameter.
These are standard results from information theory, and only the main elements of their proofs are presented here. Monotonicity is a consequence of Hölder’s inequality in the form:
@@
for $0<\alpha_{1}<\alpha_{2}$, so that $\mathbb{E}[\alpha_{1}][dc]\geq\mathbb{E}[\alpha_{2}][dc]$. The limit for $\alpha\to 0$ follows from the expansion of the cumulant generating function. For the limit $\alpha\to\infty$, Markov’s inequality:
@@
Instances of the final equality for $\bar{\mathbb{E}}=\mathbb{E}$ and $\bar{\mathbb{E}}=\mathbb{E}^{\alpha}$ generate bounds for the entropy-adjusted mean:
[…]
in the case $\alpha>0$, with the inequalities reversed in the case $\alpha<0$. The lever $\alpha$ dials the entropy-adjusted mean $\mathbb{E}[\alpha][dc]$ across the span of possible values for $dc$ between $\operatorname*{ess\,inf}[dc]$ and $\operatorname*{ess\,sup}[dc]$ and no further. Entropic risk is, in this sense, more targeted than variance adjustment as a moments-based metric of value-at-risk. The equivalent measure $\mathbb{E}^{\alpha}$ plays a central role in this result, and the gap between the mean $\mathbb{E}[dc]$ and the entropy-adjusted mean $\mathbb{E}[\alpha][dc]$ is proportional to the distance between the measure $\mathbb{E}$ and the measure $\mathbb{E}^{\alpha}$, as defined by their relative entropy.
@@ II.2 Optimising entropic risk
The investor has a choice of self-funded market strategies whose returns over the investment period are the components of the return vector $dq$. The return on the investment portfolio whose weights are the components of the weight vector $\omega$ is then $dc:=\omega\cdot dq$, the weighted sum of the returns on the market strategies.
@@
Risk aversion is controlled by the parameter $\alpha>0$, and entropic risk is monotonic in this range:
@@
$\partial\phi=\mathbb{V}^{\phi}[dq]^{-1}\mathbb{E}^{\phi}[dq]$ (28)
initialised with $\phi=0$ calibrates the unit optimal portfolio to the calibration condition when the variance is invertible. This portfolio is a property of the market return, generating the price measure $\mathbb{E}^{\phi}$ equivalent to the expectation measure $\mathbb{E}$. Applying the entropic risk metric theorem for the case $dc=\phi\cdot dq$ and $\alpha=1$ leads to the result:
$S[\bar{\mathbb{E}}|\mathbb{E}]$ $=S[\bar{\mathbb{E}}|\mathbb{E}^{\phi}]+S[\mathbb{E}^{\phi}|\mathbb{E}]-\phi\cdot(\bar{\mathbb{E}}[dq]-\mathbb{E}^{\phi}[dq])$ (29)
$\geq S[\mathbb{E}^{\phi}|\mathbb{E}]-\phi\cdot(\bar{\mathbb{E}}[dq]-\mathbb{E}^{\phi}[dq])$
Among the equivalent measures $\bar{\mathbb{E}}$ satisfying the calibration condition $\bar{\mathbb{E}}[dq]=0$, the price measure $\mathbb{E}^{\phi}$ thus has the lowest entropy relative to the expectation measure $\mathbb{E}$.
@@
where the calibration conditions in the $i$th row maximise the entropy-adjusted mean of the $i$th optimal portfolio $\phi^{i}=\phi_{1}^{i}\oplus\cdots\oplus\phi_{i}^{i}$ constituted in the lowest $i$ market tiers only. The $i$th investor then takes responsibility for determining the primary component $\psi_{i}$ of the $i$th optimal portfolio and its delta $\delta^{i}=\delta_{1}^{i}\oplus\cdots\oplus\delta_{i-1}^{i}$ to the $(i-1)$th optimal portfolio in the underlying sub-market, using the invertible linear change of variables:
@@
Stacking these deltas for $i=1,\ldots,n$ generates the optimal portfolio of the full market.
For each $i=0,\ldots,n$, define the equivalent price measure $\mathbb{E}^{i}$ with Radon-Nikodym weight:
@@
for $j=1,\ldots,i$. Starting with the expectation measure $\mathbb{E}^{0}=\mathbb{E}$, these calibrations are iterated to obtain the unit optimal portfolio $\phi^{n}$ in the full market, with its associated price measure $\mathbb{E}^{n}$ calibrated to the calibration conditions for the market strategies.
@@ III Entropic pricing and hedging
Entropic Pricing and Hedging
Fix a finite investment horizon and let $\mathbb{E}$ be the expectation measure mapping variables at the final time $T=t+dt$ to their expectations at the initial time $t$. Assume the market is tiered into a set of underlying securities with initial price vector $q$ and final price vector $Q=q+dq$ and a derivative security with initial price $p$ and final price $P=p+dp$. Consider two self-funded strategies: one comprising underlying securities only; and one that also includes the derivative security.
These two strategies offer different entropy-adjusted mean returns to the investor. When the derivative security has no intrinsic value beyond its contractual link with the underlying securities, the efficient market equilibrates to the indifferent state wherein these entropy-adjusted means align, leading to the price condition: $\mathbb{E}[\alpha][\omega\cdot dq]=\mathbb{E}[\alpha][dp+(\omega-\delta)\cdot dq]$ (42) The calibration condition is solved to derive the optimal portfolio $\omega$ of underlying securities. The hedge and price conditions are then solved simultaneously to derive the initial price $p$ from the final price $P$ of the derivative security and the hedge portfolio $\delta$ of underlying securities that re-establishes optimality for the combined portfolio.
The market of underlying securities accessible to the investor performs the two key functions of funding and hedging for the derivative security. Fix a finite investment horizon with initial time $t$ and final time $T=t+dt$, and assume the market is tiered into a set of underlying securities with initial price vector $q$ and final price vector $Q=q+dq$ and a derivative security with initial price $p$ and final price $P=p+dp$. The investor quantifies a priori the performance of the investment strategy using an initial assessment of the final economic state, encapsulated in the expectation measure $\mathbb{E}$ that maps final variables to their initial expectations. Following the strategy developed in the previous section, the optimal portfolio then maximises the entropy-adjusted mean return at a specified risk aversion $\alpha$ for the investor.
@@
where $(r+\alpha s)\,dt$ is the Lagrange multiplier used to implement the self-funding condition $p+(\omega-\delta)\cdot q=0$.
@@
The calibration, hedge and price conditions, together with the two self-funding conditions, are the foundations of entropic pricing. By avoiding the unrealistic assumptions that enable risk-neutral theory, entropic pricing can be consistently applied to pricing and hedging across a range spanning everything from algorithmic trading to XVA. These applications are explored in more depth in the following sections.
Refer to caption Refer to caption Refer to caption
Normal ($\mu_{1}=-5\%$) Normal ($\mu_{1}=0\%$) Normal ($\mu_{1}=5\%$)
Refer to caption Refer to caption Refer to caption
Binomial ($c_{1}=5\%$) Binomial ($c_{1}=10\%$) Binomial ($c_{1}=15\%$)
@@
Figure 6: Discounting is an emergent feature in entropic pricing, where the funding rate is implied as the Lagrange multiplier that implements the self-funding condition. In these studies the investment horizon is $dt=1$ and the underlying market comprises two securities with unit initial prices and independent final prices.
      In the first study, the final prices are modelled as independent normal variables with mean $q(1+\mu\,dt)$ and variance $q^{2}\sigma^{2}\,dt$. Examples are plotted for the first underlying with drift $\mu_{1}=-5\%,0\%,5\%$ and volatility $\sigma_{1}=0\%,\ldots,20\%$. The second underlying has drift $\mu_{2}=0\%$ and the implied funding rate is plotted as a function of its volatility $\sigma_{2}$. The implied funding rate is inbetween the two drifts and gravitates to the drift of the underlying with the lowest volatility.
      In the second study, the final prices are modelled as independent binomial variables with possible values $(0,1+c\,dt)$ and probabilities $(w,1-w)$. Examples are plotted for the first underlying with coupon $c_{1}=5\%,10\%,15\%$ and default probability $w_{1}=0.01\%,\ldots,10\%$. The second underlying has coupon $c_{2}=10\%$ and the implied funding rate is plotted as a function of its default probability $w_{2}$. The implied funding rate is capped by the coupons and decreases as the default probability increases.
III.1 Calibration